OpenStudy (anonymous):

A.Given the equation 5x − 4 = –2(3x + 2), solve for the variable. Explain each step and justify your process. B.Megan solved a similar equation below. Is Megan's solution correct? Explain why or why not. 3(2x − 4) = 5x − 1 6x − 12 = 5x − 1 11x − 12 = –1 11x = 11 x = 11

OpenStudy (anonymous):

ok please distribute the -2 to the parentheses on the right side

OpenStudy (anonymous):

(3x+-2)

OpenStudy (calculusxy):

\[\large 5x - 4 =-2(3x + 2) => 5x - 4 = -2(3x) + -2(2)\]

OpenStudy (calculusxy):

What is -2(3x) and -2(2) ?

OpenStudy (anonymous):

^^ would that be the top piece for my answer cuz i have to put the steps on how i got the answer

OpenStudy (calculusxy):

Yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

just finish going through how to get it and you will know what to put for your answer but don't worry about it yet

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what next

OpenStudy (calculusxy):

What did you get after your distributed -2 to the numbers inside of the parenthesis?

OpenStudy (anonymous):

-2(3x)=-6 -2(2)= -4

OpenStudy (calculusxy):

Since you multiplied -2 with 3x, it's -6x. \[\large 5x - 4 = -6x -4 => 5x - 4 + 6x = -6x - 4 + 6x => 11x -4 = -4\]

OpenStudy (calculusxy):

\[11x -4 = -4\]

OpenStudy (anonymous):

thats it?

OpenStudy (calculusxy):

And then solve for x.

OpenStudy (anonymous):

like that

OpenStudy (anonymous):

so far

OpenStudy (anonymous):

its worth 10 points i really need to pass this question lol

OpenStudy (anonymous):

anyone??

OpenStudy (calculusxy):

Instead of putting this symbol \[=>\] just press 'enter' and go to the line below. also you need to solve for x. \[\large 11x - 4 = -4 \] \[\large 11x - 4 + 4 = -4 + 4\] \[\large 11x = 0\] this shows that x is 0

OpenStudy (anonymous):

so that

OpenStudy (anonymous):

what about part B